A hot line with a full free number receives an average of 4 calls per hour for any given hour.  Find the probability that it will receive exactly 5 calls



1
Expert's answer
2021-06-01T13:10:17-0400

"P(X=x) = \frac{e^{-λ}λ^x}{x!} \\

λ=4 \\

P(X=5) = \frac{e^{-4}4^5}{5!} = 0.1563"